# Ergodicity, KAM, FPUT

### v1

@@ 2 KAM
  • As a first example, a condition for the existence of the shift $\bf a$}, in the first of Eq.(2.6), is $\varepsilon_{0}$ small: if $J_{0,-}^{-1}\varepsilon_{0}\rho_{0}^{-1}<\frac{1}{\gamma_{0}}$ and $\gamma_{0}$ is large enough, the equation ${\bf a}=-J_{0}^{-1}{\mbox{\boldmath$\partial$}}_{\bf a}{\overline{f}}({\bf a})$ has a solution $\bf a$} with $|{\bf a}|<J_{0,-}^{-1}\varepsilon_{0}<\frac{1}{\gamma_{0}}\rho_{0}$: see Appendix A  where the choice $\gamma_{0}=3N$ is proposed.
However the $\Phi$ in Eq.(2.6) does not make sense (in general); instead it will “be approximated”, and properly defined, by $\Phi_{0}({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}})$ as:
${{}-\kern-8.0pt\sum_{{\bf 0}\neq{\mbox{\boldmath$\nu$}}\in\mathcal{Z}^{N}}\kern-5.0pt\frac{f_{0,{\mbox{\boldmath$\nu$}}}({\bf A}^{\prime}+{\bf a})e^{i{\mbox{\boldmath$\alpha$}}\cdot{\mbox{\boldmath$\nu$}}}}{i{\mbox{\boldmath$\omega$}}_{0}\cdot{\mbox{\boldmath$\nu$}}}{\Big{(}1-\frac{J_{0}({\bf A}^{\prime}+{\bf a})\cdot{\mbox{\boldmath$\nu$}}}{{\mbox{\boldmath$\omega$}}_{0}\cdot{\mbox{\boldmath$\nu$}}}+\Big{(}\frac{J_{0}({\bf A}^{\prime}+{\bf a})\cdot{\mbox{\boldmath$\nu$}}}{{\mbox{\boldmath$\omega$}}_{0}\cdot{\mbox{\boldmath$\nu$}}}\Big{)}^{2}\Big{)}}}$
From $\varepsilon_{0}=\|{\partial}_{{{\bf A}}^{\prime}}f_{0}\|_{\rho_{0},\kappa_{0}}+\frac{1}{\rho_{0}}\|{\partial}_{{\mbox{\boldmath$\alpha$}}}f_{0}\|_{\rho_{0},\kappa_{0}}$ the dimensional inequalities (Eq.2.7):
$|{\mbox{\boldmath$\partial$}}_{{{\bf A}}^{\prime}}f_{0,{\mbox{\boldmath$\nu$}}}({{\bf A}}^{\prime}+{\bf a})|\leq\varepsilon_{0}e^{-\kappa_{0}|{\mbox{\boldmath$\nu$}}|},\qquad\rho_{0}|{\mbox{\boldmath$\nu$}}||f_{0,{\mbox{\boldmath$\nu$}}}({{\bf A}}^{\prime}+{\bf a})|<\varepsilon_{0}e^{-\kappa_{0}\cdot|{\mbox{\boldmath$\nu$}}|}$ (2.9)
@@
${{}\|\Phi_{0}\|_{\frac{2}{3}\rho_{0},\kappa_{0}}\leq\varepsilon_{0}\rho_{0}\kern-5.69054pt\sum_{{\mbox{\boldmath$\nu$}}\neq{\bf 0}}\frac{|{\mbox{\boldmath$\nu$}}|e^{-\kappa_{0}|{\mbox{\boldmath$\nu$}}|}}{|{\mbox{\boldmath$\omega$}}_{0}\cdot{\mbox{\boldmath$\nu$}}|}\Big{(}1\kern-5.69054pt+\frac{|J_{0}{\mbox{\boldmath$\nu$}}|\rho_{0}}{|{\mbox{\boldmath$\omega$}}_{0}\cdot{\mbox{\boldmath$\nu$}}|}+\kern-5.69054pt\Big{(}\frac{|J_{0}{\mbox{\boldmath$\nu$}}|\rho_{0}}{|{\mbox{\boldmath$\omega$}}_{0}\cdot{\mbox{\boldmath$\nu$}}|}\Big{)}^{2}\Big{)}\leq\gamma_{1}\frac{\varepsilon_{0}}{C}\,\kappa_{0}^{-c_{1}}}$
@@
${{}\leq\widetilde{\gamma}\varepsilon_{0}\sum_{{\mbox{\boldmath$\nu$}}\neq{\bf 0}}\frac{|{\mbox{\boldmath$\nu$}}|^{2}e^{-\kappa_{0}|{\mbox{\boldmath$\nu$}}|}}{|{\mbox{\boldmath$\omega$}}_{0}\cdot{\mbox{\boldmath$\nu$}}|}\Big{(}1+\frac{|J_{0}{\mbox{\boldmath$\nu$}}|\rho_{0}}{|{\mbox{\boldmath$\omega$}}_{0}\cdot{\mbox{\boldmath$\nu$}}|}+\Big{(}\frac{|J_{0}{\mbox{\boldmath$\nu$}}|\rho_{0}}{|{\mbox{\boldmath$\omega$}}_{0}\cdot{\mbox{\boldmath$\nu$}}|}\Big{)}^{2}\Big{)}\leq\widehat{\gamma}\frac{\varepsilon_{0}}{C}\,\kappa_{0}^{-\widehat{c}}}$
with $\widetilde{\gamma},\widehat{\gamma},\widehat{c}$ dimensionless constants.  [^9] E.g. $\widehat{c}=4N+2$, $\widehat{g}=\widetilde{\gamma}\sum_{\mbox{\boldmath$\nu$}}|{\mbox{\boldmath$\nu$}}|^{4N+2}e^{-\kappa_{0}\,|{\mbox{\boldmath$\nu$}}|}$, $\widetilde{\gamma}=7$. Remark that the bound on the second term in the first line of Eq.2.12 does not involve a change in $\kappa_{0}$, unlike the first term (where $\rho$ is reduced).
@@
The ’new’ perturbation will be, by Eq.(2.5), $f_{1}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})=\sum_{j=1}^{3}F_{j}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})$ if expressed in the mixed variables $({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})$. Hence it will eventually be necessary to express $\alpha$ via $({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})$ by inverting the second in Eq.2.4 as ${\mbox{\boldmath$\alpha$}}={\mbox{\boldmath$\alpha$}}^{\prime}+\Delta({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})$: so that the perturbation acting in $H_{1}$ will be, properly, $f_{1}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})=\sum_{j=1}^{3}F_{j}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime}+\Delta({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime}))$.
To solve for the implicit function $\Delta({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})$ in the second of Eq.2.4 , an argument can be used similar to the one for the solution of ${\bf a}=-J_{0}^{-1}{\mbox{\boldmath$\partial$}}_{\bf a}{\overline{f}}_{0}({\bf a})$ in Appendix A . In Appendix B the similar analysis is adopted to construct $\Delta({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})$ and a proof that it can be defined on ${\mathcal{C}}(\frac{1}{2}\rho_{0},\kappa_{0}-2\delta_{0})$ subject to the bound:
@@
for dimensionless ${\overline{\gamma}},{\overline{c}}$, and for $\delta_{0}>0$ to be determined and depending on $\varepsilon_{0},\kappa_{0}$. However it is convenient to postpone the study of $\Delta$ and to obtain first bounds on $f^{\prime}_{1}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})=\sum_{j=1}^{3}F_{j}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})$ in a domain ${\mathcal{C}}(\widetilde{\rho},\kappa_{0})$. The conditions for the definitions of $F_{j}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})$, i.e. for the definition of the generating function $\Phi_{0}$ to be defined in the mixed variables $({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})$, are summarized in Eq.2.10
Under such conditions the $F_{j}$ composing $f^{\prime}_{1}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})$, in the lines labeled $(j),j>0$ in Eq.(2.5), can be easily bounded dimensionally as follows.
  • (a) $F_{0}$ is $\frac{1}{2}(J_{0}{{\bf A}}^{\prime}\cdot{{\bf A}}^{\prime})+{\mbox{\boldmath$\omega$}}_{0}\cdot{{\bf A}}^{\prime}$ for ${{\bf A}}^{\prime}\in C(\frac{1}{2}\rho_{0})$.
@@
    ${{}|{\partial}_{{{\bf A}}^{\prime}}F_{2}|_{\widetilde{\rho},\kappa_{0}}+\frac{1}{\widetilde{\rho}}|{\partial}_{\mbox{\boldmath$\alpha$}}F_{2}|_{\widetilde{\rho},\kappa_{0}}\leq\gamma_{3}\varepsilon_{0}\frac{\rho_{0}}{\widetilde{\rho}}\frac{\varepsilon_{0}}{C}\kappa_{0}^{-c_{3}}}$
@@
    ${{}F_{3}({\bf A}^{\prime},{\bf z}){\buildrel def\over{=}}({\mbox{\boldmath$\omega$}}_{0}\kern-2.84526pt+\kern-2.84526ptJ_{0}({\bf A}^{\prime}\kern-2.84526pt+\kern-2.84526pt{\bf a})\cdot{\mbox{\boldmath$\partial$}}_{\mbox{\boldmath$\alpha$}}\Phi_{0}\kern-2.84526pt+\kern-2.84526ptf_{0}({\bf A}^{\prime}\kern-2.84526pt+\kern-2.84526pt{\bf a},{\mbox{\boldmath$\alpha$}})\kern-2.84526pt-\kern-2.84526pt{\overline{f}}_{0}({\bf A}^{\prime}+{\bf a}))}$
    ${{}=-\sum_{{\bf 0}\neq{\mbox{\boldmath$\nu$}}}f_{0,{\mbox{\boldmath$\nu$}}}({\bf A}^{\prime}+{\bf a})\frac{(J_{0}({\bf A}^{\prime}+{\bf a})\cdot{\mbox{\boldmath$\nu$}})^{3}}{({\mbox{\boldmath$\omega$}}_{0}\cdot{\mbox{\boldmath$\nu$}})^{3}}e^{i{\mbox{\boldmath$\alpha$}}\cdot{\mbox{\boldmath$\nu$}}}\leq\widetilde{\gamma}\frac{\varepsilon_{0}}{\rho_{0}}\widetilde{\rho}^{3}\kappa_{0}^{-c}}$
    and can be bounded in ${\mathcal{C}}_{\widetilde{\rho},\kappa_{0}}$ by $\widetilde{\gamma}\frac{\varepsilon_{0}}{\rho_{0}}\widetilde{\rho}^{3}\kappa_{0}^{-c}$ for suitable dimendionless $\widetilde{\gamma},c$, and in the variables $({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})\in{\mathcal{C}}_{\widetilde{\rho},\kappa_{0}}$ as in c) above:
    ${{}|{\partial}_{{{\bf A}}^{\prime}}F_{3}|_{\widetilde{\rho},\kappa_{0}}+\frac{1}{\widetilde{\rho}}|{\partial}_{\mbox{\boldmath$\alpha$}}F_{3}|_{\widetilde{\rho},\kappa_{0}}\leq\gamma_{4}\varepsilon_{0}\Big{(}\frac{\widetilde{\rho}}{\rho_{0}}\Big{)}^{2}\kappa_{0}^{-c_{4}}}$
@@
${{}J_{1}=J_{0},\qquad|{\mbox{\boldmath$\partial$}}_{{\bf A}^{\prime}}F_{1}({\bf A}^{\prime})|\leq{\overline{\gamma}}\varepsilon_{0}\Big{(}\frac{\widetilde{\rho}}{\rho_{0}}\Big{)},}$
${{}|{\mbox{\boldmath$\partial$}}_{{\bf A}^{\prime}}F_{2}({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}})|+\frac{1}{\widetilde{\rho}}|{\mbox{\boldmath$\partial$}}_{\mbox{\boldmath$\alpha$}}F_{2}({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}})|\leq{\overline{\gamma}}\varepsilon_{0}\frac{\varepsilon_{0}}{C}\frac{\rho_{0}}{\widetilde{\rho}}\kappa_{0}^{-{\overline{c}}},}$
${{}|{\mbox{\boldmath$\partial$}}_{{\bf A}^{\prime}}F_{3}({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}})|+\frac{1}{\widetilde{\rho}}|{\mbox{\boldmath$\partial$}}_{\mbox{\boldmath$\alpha$}}F_{3}({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}})|\leq{\overline{\gamma}}\varepsilon_{0}\Big{(}\frac{\widetilde{\rho}}{\rho_{0}}\Big{)}^{2}\kappa_{0}^{-{\overline{c}}}}$
for ${\overline{\gamma}},{\overline{c}}$ constants (dimensionless), $({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}})\in{\mathcal{C}}(\widetilde{\rho},\kappa_{0})$.
@@
Thus the corresponding $F^{\prime}_{j}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime}){\buildrel def\over{=}}F_{j}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})$ admit the same bounds as Eq.2.18, but in the smaller domain ${\mathcal{C}}(\frac{1}{2}\rho_{0},\kappa_{0}-2\delta_{0})$ with $\delta_{0}\leq{\overline{\gamma}}(\frac{\varepsilon_{0}}{C})\kappa_{0}^{-{\overline{c}}}<\frac{1}{2}\kappa_{0}$, by simply suitably increasing the constants ${\overline{\gamma}},{\overline{c}}$, as discussed in AppendixB .
@@
Iterate the renormalization transformation $({{\bf A}}_{n-1},{\mbox{\boldmath$\alpha$}}_{n-1})\to({{\bf A}}^{\prime}={{\bf A}}_{n},{\mbox{\boldmath$\alpha$}}^{\prime}={\mbox{\boldmath$\alpha$}}_{n})$ transforming the Hamiltonian $H_{n-1}({{\bf A}}_{n-1},{\mbox{\boldmath$\alpha$}}_{n-1})$, on ${\mathcal{C}}(\rho_{n-1},\kappa_{n-1})$ with pertubation $f_{n-1}$, into $H_{n}({{\bf A}}_{n},{\mbox{\boldmath$\alpha$}}_{n})$, on ${\mathcal{C}}(\rho_{n},\kappa_{n})$ with pertubation $f_{n}$. Suppose inductively, for all $n\geq 0$:
@@
Then, for $({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})\in{\mathcal{C}}(\rho_{n},\kappa_{n})$, the $\varepsilon_{n}=\|{\mbox{\boldmath$\partial$}}_{{\bf A}^{\prime}}f_{n}\|_{\rho_{n},\kappa_{n}}+\frac{1}{\rho_{n}}\|{\mbox{\boldmath$\partial$}}_{{\mbox{\boldmath$\alpha$}}}f_{n}\|_{\rho_{n},\kappa_{n}}$ are bounded, if $\kappa_{n}>\frac{1}{2}\kappa_{0}$, by:
@@
The recursion, Eq. 2.19 , for $\frac{\varepsilon_{n}}{C}$ can be iterated for $n\geq 1$, as long as the conditions Eq.2.10 remain valid, finding $\frac{\varepsilon_{n}}{C}\leq\Lambda(\frac{\varepsilon_{n-1}}{C})^{\xi}$ for $\Lambda={\overline{\gamma}}(\frac{\kappa_{0}}{2})^{-{\overline{c}}}$ and $\gamma,c,\xi=\frac{3}{2}$ dimensionless constants. If $\eta_{n}{\buildrel def\over{=}}\sigma\frac{\varepsilon_{n}}{C}$ with $\sigma=\Lambda^{2}$ this becomes simply $\eta_{n}\leq\eta_{0}^{\xi^{n-1}}$, and consequently $\rho_{n}\leq\Lambda^{-\frac{n(n+1)}{2}}\eta_{0}^{\xi^{n}-1}\rho_{0}=O(\eta_{0}^{\xi^{n}})\rho_{0}=O((\Lambda^{2}\frac{\varepsilon_{0}}{C})^{\xi^{n}})\rho_{0}$.

### v2

@@ 2 KAM
  • As a first example, a condition for the existence of the shift $\bf a$}, in the first of Eq.(2.6), is $\varepsilon_{0}$ small: if $J_{0,-}^{-1}\varepsilon_{0}\rho_{0}^{-1}<\frac{1}{\gamma_{0}}$ and $\gamma_{0}$ is large enough, the equation ${\bf a}=-J_{0}^{-1}{\mbox{$\partial$}}_{\bf a}{\overline{f}}({\bf a})$ has a solution $\bf a$} with $|{\bf a}|<J_{0,-}^{-1}\varepsilon_{0}<\frac{1}{\gamma_{0}}\rho_{0}$: see Appendix A  where the choice $\gamma_{0}=3N$ is proposed.
However the $\Phi$ in Eq.(2.6) does not make sense (in general); instead it will “be approximated”, and properly defined, by $\Phi_{0}({\bf A}^{\prime},{\mbox{$\alpha$}})$ as:
${{}-\kern-8.0pt\sum_{{\bf 0}\neq{\mbox{$\nu$}}\in\mathcal{Z}^{N}}\kern-5.0pt\frac{f_{0,{\mbox{$\nu$}}}({\bf A}^{\prime}+{\bf a})e^{i{\mbox{$\alpha$}}\cdot{\mbox{$\nu$}}}}{i{\mbox{$\omega$}}_{0}\cdot{\mbox{$\nu$}}}{\Big(1-\frac{J_{0}({\bf A}^{\prime}+{\bf a})\cdot{\mbox{$\nu$}}}{{\mbox{$\omega$}}_{0}\cdot{\mbox{$\nu$}}}+\Big(\frac{J_{0}({\bf A}^{\prime}+{\bf a})\cdot{\mbox{$\nu$}}}{{\mbox{$\omega$}}_{0}\cdot{\mbox{$\nu$}}}\Big)^{2}\Big)}}$
From $\varepsilon_{0}=\|{\partial}_{{{\bf A}}^{\prime}}f_{0}\|_{\rho_{0},\kappa_{0}}+\frac{1}{\rho_{0}}\|{\partial}_{{\mbox{$\alpha$}}}f_{0}\|_{\rho_{0},\kappa_{0}}$ the dimensional inequalities (Eq.2.7):
$|{\mbox{$\partial$}}_{{{\bf A}}^{\prime}}f_{0,{\mbox{$\nu$}}}({{\bf A}}^{\prime}+{\bf a})|\leq\varepsilon_{0}e^{-\kappa_{0}|{\mbox{$\nu$}}|},\qquad\rho_{0}|{\mbox{$\nu$}}||f_{0,{\mbox{$\nu$}}}({{\bf A}}^{\prime}+{\bf a})|<\varepsilon_{0}e^{-\kappa_{0}\cdot|{\mbox{$\nu$}}|}$ (2.9)
@@
${{}\|\Phi_{0}\|_{\frac{2}{3}\rho_{0},\kappa_{0}}\leq\varepsilon_{0}\rho_{0}\kern-5.69054pt\sum_{{\mbox{$\nu$}}\neq{\bf 0}}\frac{|{\mbox{$\nu$}}|e^{-\kappa_{0}|{\mbox{$\nu$}}|}}{|{\mbox{$\omega$}}_{0}\cdot{\mbox{$\nu$}}|}\Big(1\kern-5.69054pt+\frac{|J_{0}{\mbox{$\nu$}}|\rho_{0}}{|{\mbox{$\omega$}}_{0}\cdot{\mbox{$\nu$}}|}+\kern-5.69054pt\Big(\frac{|J_{0}{\mbox{$\nu$}}|\rho_{0}}{|{\mbox{$\omega$}}_{0}\cdot{\mbox{$\nu$}}|}\Big)^{2}\Big)\leq\gamma_{1}\frac{\varepsilon_{0}}{C}\,\kappa_{0}^{-c_{1}}}$
@@
${{}\leq\widetilde{\gamma}\varepsilon_{0}\sum_{{\mbox{$\nu$}}\neq{\bf 0}}\frac{|{\mbox{$\nu$}}|^{2}e^{-\kappa_{0}|{\mbox{$\nu$}}|}}{|{\mbox{$\omega$}}_{0}\cdot{\mbox{$\nu$}}|}\Big(1+\frac{|J_{0}{\mbox{$\nu$}}|\rho_{0}}{|{\mbox{$\omega$}}_{0}\cdot{\mbox{$\nu$}}|}+\Big(\frac{|J_{0}{\mbox{$\nu$}}|\rho_{0}}{|{\mbox{$\omega$}}_{0}\cdot{\mbox{$\nu$}}|}\Big)^{2}\Big)\leq\widehat{\gamma}\frac{\varepsilon_{0}}{C}\,\kappa_{0}^{-\widehat{c}}}$
with $\widetilde{\gamma},\widehat{\gamma},\widehat{c}$ dimensionless constants.  [^9] E.g. $\widehat{c}=4N+2$, $\widehat{g}=\widetilde{\gamma}\sum_{\mbox{$\nu$}}|{\mbox{$\nu$}}|^{4N+2}e^{-\kappa_{0}\,|{\mbox{$\nu$}}|}$, $\widetilde{\gamma}=7$. Remark that the bound on the second term in the first line of Eq.2.12 does not involve a change in $\kappa_{0}$, unlike the first term (where $\rho$ is reduced).
@@
The ’new’ perturbation will be, by Eq.(2.5), $f_{1}({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime})=\sum_{j=1}^{3}F_{j}({{\bf A}}^{\prime},{\mbox{$\alpha$}})$ if expressed in the mixed variables $({{\bf A}}^{\prime},{\mbox{$\alpha$}})$. Hence it will eventually be necessary to express $\alpha$ via $({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime})$ by inverting the second in Eq.2.4 as ${\mbox{$\alpha$}}={\mbox{$\alpha$}}^{\prime}+\Delta({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime})$: so that the perturbation acting in $H_{1}$ will be, properly, $f_{1}({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime})=\sum_{j=1}^{3}F_{j}({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime}+\Delta({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime}))$.
To solve for the implicit function $\Delta({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime})$ in the second of Eq.2.4 , an argument can be used similar to the one for the solution of ${\bf a}=-J_{0}^{-1}{\mbox{$\partial$}}_{\bf a}{\overline{f}}_{0}({\bf a})$ in Appendix A . In Appendix B the similar analysis is adopted to construct $\Delta({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime})$ and a proof that it can be defined on ${\mathcal{C}}(\frac{1}{2}\rho_{0},\kappa_{0}-2\delta_{0})$ subject to the bound:
@@
for dimensionless ${\overline{\gamma}},{\overline{c}}$, and for $\delta_{0}>0$ to be determined and depending on $\varepsilon_{0},\kappa_{0}$.
However it is convenient to postpone the study of $\Delta$ and to obtain first bounds on $f^{\prime}_{1}({{\bf A}}^{\prime},{\mbox{$\alpha$}})=\sum_{j=1}^{3}F_{j}({{\bf A}}^{\prime},{\mbox{$\alpha$}})$ in a domain ${\mathcal{C}}(\widetilde{\rho},\kappa_{0})$.
The conditions for the definitions of $F_{j}({{\bf A}}^{\prime},{\mbox{$\alpha$}})$, i.e. for the definition of the generating function $\Phi_{0}$ to be defined in the mixed variables $({{\bf A}}^{\prime},{\mbox{$\alpha$}})$, are summarized in Eq.2.10
Under such conditions the $F_{j}$ composing $f^{\prime}_{1}({{\bf A}}^{\prime},{\mbox{$\alpha$}})$, in the lines labeled $(j),j>0$ in Eq.(2.5), can be easily bounded dimensionally as follows.
  • (a) $F_{0}$ is $\frac{1}{2}(J_{0}{{\bf A}}^{\prime}\cdot{{\bf A}}^{\prime})+{\mbox{$\omega$}}_{0}\cdot{{\bf A}}^{\prime}$ for ${{\bf A}}^{\prime}\in C(\frac{1}{2}\rho_{0})$.
@@
    ${{}|{\partial}_{{{\bf A}}^{\prime}}F_{2}|_{\widetilde{\rho},\kappa_{0}}+\frac{1}{\widetilde{\rho}}|{\partial}_{\mbox{$\alpha$}}F_{2}|_{\widetilde{\rho},\kappa_{0}}\leq\gamma_{3}\varepsilon_{0}\frac{\rho_{0}}{\widetilde{\rho}}\frac{\varepsilon_{0}}{C}\kappa_{0}^{-c_{3}}}$
@@
    ${{}F_{3}({\bf A}^{\prime},{\bf z}){\mathrel{\mathop{\kern 0.0pt=}\limits^{def}}}({\mbox{$\omega$}}_{0}\kern-2.84526pt+\kern-2.84526ptJ_{0}({\bf A}^{\prime}\kern-2.84526pt+\kern-2.84526pt{\bf a})\cdot{\mbox{$\partial$}}_{\mbox{$\alpha$}}\Phi_{0}\kern-2.84526pt+\kern-2.84526ptf_{0}({\bf A}^{\prime}\kern-2.84526pt+\kern-2.84526pt{\bf a},{\mbox{$\alpha$}})\kern-2.84526pt-\kern-2.84526pt{\overline{f}}_{0}({\bf A}^{\prime}+{\bf a}))}$
    ${{}=-\sum_{{\bf 0}\neq{\mbox{$\nu$}}}f_{0,{\mbox{$\nu$}}}({\bf A}^{\prime}+{\bf a})\frac{(J_{0}({\bf A}^{\prime}+{\bf a})\cdot{\mbox{$\nu$}})^{3}}{({\mbox{$\omega$}}_{0}\cdot{\mbox{$\nu$}})^{3}}e^{i{\mbox{$\alpha$}}\cdot{\mbox{$\nu$}}}\leq\widetilde{\gamma}\frac{\varepsilon_{0}}{\rho_{0}}\widetilde{\rho}^{3}\kappa_{0}^{-c}}$
    and can be bounded in ${\mathcal{C}}_{\widetilde{\rho},\kappa_{0}}$ by $\widetilde{\gamma}\frac{\varepsilon_{0}}{\rho_{0}}\widetilde{\rho}^{3}\kappa_{0}^{-c}$ for suitable dimendionless $\widetilde{\gamma},c$, and in the variables $({{\bf A}}^{\prime},{\mbox{$\alpha$}})\in{\mathcal{C}}_{\widetilde{\rho},\kappa_{0}}$ as in c) above:
    ${{}|{\partial}_{{{\bf A}}^{\prime}}F_{3}|_{\widetilde{\rho},\kappa_{0}}+\frac{1}{\widetilde{\rho}}|{\partial}_{\mbox{$\alpha$}}F_{3}|_{\widetilde{\rho},\kappa_{0}}\leq\gamma_{4}\varepsilon_{0}\Big(\frac{\widetilde{\rho}}{\rho_{0}}\Big)^{2}\kappa_{0}^{-c_{4}}}$
@@
${{}J_{1}=J_{0},\qquad|{\mbox{$\partial$}}_{{\bf A}^{\prime}}F_{1}({\bf A}^{\prime})|\leq{\overline{\gamma}}\varepsilon_{0}\Big(\frac{\widetilde{\rho}}{\rho_{0}}\Big),}$
${{}|{\mbox{$\partial$}}_{{\bf A}^{\prime}}F_{2}({\bf A}^{\prime},{\mbox{$\alpha$}})|+\frac{1}{\widetilde{\rho}}|{\mbox{$\partial$}}_{\mbox{$\alpha$}}F_{2}({\bf A}^{\prime},{\mbox{$\alpha$}})|\leq{\overline{\gamma}}\varepsilon_{0}\frac{\varepsilon_{0}}{C}\frac{\rho_{0}}{\widetilde{\rho}}\kappa_{0}^{-{\overline{c}}},}$
${{}|{\mbox{$\partial$}}_{{\bf A}^{\prime}}F_{3}({\bf A}^{\prime},{\mbox{$\alpha$}})|+\frac{1}{\widetilde{\rho}}|{\mbox{$\partial$}}_{\mbox{$\alpha$}}F_{3}({\bf A}^{\prime},{\mbox{$\alpha$}})|\leq{\overline{\gamma}}\varepsilon_{0}\Big(\frac{\widetilde{\rho}}{\rho_{0}}\Big)^{2}\kappa_{0}^{-{\overline{c}}}}$
for ${\overline{\gamma}},{\overline{c}}$ constants (dimensionless), $({\bf A}^{\prime},{\mbox{$\alpha$}})\in{\mathcal{C}}(\widetilde{\rho},\kappa_{0})$.
@@
Thus the corresponding $F^{\prime}_{j}({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime}){\mathrel{\mathop{\kern 0.0pt=}\limits^{def}}}F_{j}({{\bf A}}^{\prime},{\mbox{$\alpha$}})$ admit the same bounds as Eq.2.18, but in the smaller domain ${\mathcal{C}}(\frac{1}{2}\rho_{0},\kappa_{0}-2\delta_{0})$ with $\delta_{0}\leq{\overline{\gamma}}(\frac{\varepsilon_{0}}{C})\kappa_{0}^{-{\overline{c}}}<\frac{1}{2}\kappa_{0}$, by simply suitably increasing the constants ${\overline{\gamma}},{\overline{c}}$, as discussed in AppendixB .
@@
Iterate the renormalization transformation $({{\bf A}}_{n-1},{\mbox{$\alpha$}}_{n-1})\to({{\bf A}}^{\prime}={{\bf A}}_{n},{\mbox{$\alpha$}}^{\prime}={\mbox{$\alpha$}}_{n})$ transforming the Hamiltonian $H_{n-1}({{\bf A}}_{n-1},{\mbox{$\alpha$}}_{n-1})$, on ${\mathcal{C}}(\rho_{n-1},\kappa_{n-1})$ with pertubation $f_{n-1}$, into $H_{n}({{\bf A}}_{n},{\mbox{$\alpha$}}_{n})$, on ${\mathcal{C}}(\rho_{n},\kappa_{n})$ with pertubation $f_{n}$. Suppose inductively, for all $n\geq 0$:
@@
Then, for $({\bf A}^{\prime},{\mbox{$\alpha$}}^{\prime})\in{\mathcal{C}}(\rho_{n},\kappa_{n})$, the $\varepsilon_{n}=\|{\mbox{$\partial$}}_{{\bf A}^{\prime}}f_{n}\|_{\rho_{n},\kappa_{n}}+\frac{1}{\rho_{n}}\|{\mbox{$\partial$}}_{{\mbox{$\alpha$}}}f_{n}\|_{\rho_{n},\kappa_{n}}$ are bounded, if $\kappa_{n}>\frac{1}{2}\kappa_{0}$, by:
@@
The recursion, Eq. 2.19 , for $\frac{\varepsilon_{n}}{C}$ can be iterated for $n\geq 1$, as long as the conditions Eq.2.10 remain valid, finding $\frac{\varepsilon_{n}}{C}\leq\Lambda(\frac{\varepsilon_{n-1}}{C})^{\xi}$ for $\Lambda={\overline{\gamma}}(\frac{\kappa_{0}}{2})^{-{\overline{c}}}$ and $\gamma,c,\xi=\frac{3}{2}$ dimensionless constants. If $\eta_{n}{\mathrel{\mathop{\kern 0.0pt=}\limits^{def}}}\sigma\frac{\varepsilon_{n}}{C}$ with $\sigma=\Lambda^{2}$ this becomes simply $\eta_{n}\leq\eta_{0}^{\xi^{n-1}}$, and consequently $\rho_{n}\leq\Lambda^{-\frac{n(n+1)}{2}}\eta_{0}^{\xi^{n}-1}\rho_{0}=O(\eta_{0}^{\xi^{n}})\rho_{0}=O((\Lambda^{2}\frac{\varepsilon_{0}}{C})^{\xi^{n}})\rho_{0}$.